# Tensor network methods for quantum-inspired image processing and classical optics

### v1

@@ III Matrix Product States and Tensor Tree Networks
Nevertheless, it is clear that the error goes to zero rather differently due to the maximal number of parameters allowed by the SVD in these kinds of networks. We can show that a large $n$, maximally random tensor tree has at most $2^{n}$ parameters. We take the example of the boundary binary tree (see right Fig. 7), the maximal number of parameter is $D_{\mathrm{max}}(n)=n(2\times 2)+\sum_{k=1}^{\log_{2}(n)-2}\frac{n}{2^{k}}\left(2^{2^{k+1}-2}\times 2^{2^{k+1}-1}\right)+2\times 2^{n-2}$. The first term $n(2\times 2)$ corresponds to the $n$ outside branch tensors and the last term corresponds to the two central tensors. The largest term inside the sum is $2^{n}-1$ so $D_{\mathrm{max}}(n)=4n+2^{n-1}+2.2^{n-2}=2^{n}+4n$, implying a value (up to a tiny $\mathcal{O}(n)$ term) of $\xi_{\infty}^{\mathrm{TTN}}=1$ for large $n$. A similar counting operation can be done for bulk tensor trees, and would have lead to the same result. It means that in the thermodynamic limit, in contrast to matrix product states, $\mathrm{TTN}$ always act as a compressor ($\xi\leq 1$) regardless of the level of approximation which is already an interesting fact on its own. Therefore, tensor trees offer greater compressive power in the intermediate regime (see Fig. 8), as the number of parameters approaches the compression limit with increasing system size.
This is illustrated in the table below, where tensor trees clearly outperform MPS and continue to compress the image even for very small reconstruction error, whereas MPS fail to do so.
@@
Table 1: Compression ratio for several errors with $L=2048$. The three firsts tensor nets use FRQI with respectively a random walk (SARW), snake-like (S) and a Hilbert curve (H) path encoding. The fourth is the quantics tensor train and the last is a quantics encoding on a bulk tensor tree. Tensor trees manage to still be compressive (last column) even for very small errors, whereas MPS do not.
We expect this behavior to be generic in average and not specific to this particular image. Although comparisons with standard image compression algorithms are not showed here, we mention that tensor networks might have a big advantage at very large size and relative high error. A comprehensive benchmark—although beyond the scope of this paper—on the performance of tensor networks across large datasets of diverse images would be necessary to rigorously demonstrate their relative strengths and weaknesses. The last section of this work will be devoted to optical applications that could be significantly sped up using this formalism.
IV Image processing and image formation
@@
Figure 9: Operations on images as MPS/MPO contraction on the left and TTN/TTO contraction on the right. The takeaway message is that theses contractions can be done very efficiently.
@@
Equivalently, it can be written as a unitary operator $\hat{U}$ on the space of qubits such that [^6] In this paper, we choose the notation quantum computation notation $\{\hat{X},\hat{Y},\hat{Z}\}$ instead of $\{\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z}\}$ for the Pauli matrices.
@@
In the following, we shall see how to incorporate the previous results into a realistic optical system. The simplest of calculations is to compute the point spread function of a classical optical system that contains phase aberrations. A monochromatic coherent field can be expressed as the plane wave
$E(r,\theta)=A(r,\theta)\mathrm{e}^{-ik\Phi(r,\theta)},$ (7)
@@
$E_{\mathrm{f}}(u,v)=\int r\mathrm{d}r\int\mathrm{d}\theta E(r,\theta)\mathrm{e}^{-ikr(u\cos\theta+v\sin\theta)},$ (8)
where $(u,v)$ are spatial coordinates in the focal plane, and $A(r,\theta)$ is the pupil function, typically a circular aperture. The point spread function is simply given by
$\text{PSF}(u,v)=\left|E_{\mathrm{f}}(u,v)\right|^{2},$ (9)
which is the intensity of the field at the focal plane.
Figure 10: Three classic apertures considered for the simulations plus a continuous phase field living on them. The middle case will be surprisingly easy to approximate with MPS while the others will turn up to be more difficult.
The classical Fraunhofer propagation of the continuous field $E_{0}$ can be seen as a unitary propagation of a non-normalized quantum state $|\phi\rangle$ living in the Hilbert space of infinitely many qubits. The non-unitary operation that computes the PSF of the optical system is, in this context, defined by the non-unitary process of quantum measurement. Mathematically, it boils down to the following dictionnary
$E_{0}\leftrightarrow|\phi\rangle,$ (10)
$\mathcal{F}\{\bullet\}\leftrightarrow\hat{U}_{\mathrm{qft}}=\hat{U}^{(x)}_{\mathrm{qft}}\otimes\hat{U}^{(y)}_{\mathrm{qft}}$, (11)
$E_{\mathrm{f}}\leftrightarrow|\psi\rangle=\hat{U}_{\mathrm{qft}}|\phi\rangle,$ (12)
$\mathrm{PSF}\leftrightarrow|\mathrm{PSF}\rangle=\frac{|\psi\rangle|\psi\rangle^{\dagger}}{|\langle\psi|\psi\rangle|^{2}}.$ (13)
As said earlier, a low-rank matrix product operator construction of the QFT has recently been developed and could be implemented here. To take advantage of this MPO formulation, the field $|\phi\rangle$ requires also a low-rank form. A summary of the calculation steps is illustrated in Fig. 11.
@@
Contrary to the standard PSF calculation, the matrix product state formulation allows photon sampling (see Fig. 11) without ever computing the photon statistical distribution, ie. the PSF. The perfect sampling algorithm can be used to sample from this distribution Stoudenmire and White (2010); Ferris and Vidal (2012). As the name suggests, the algorithm draws perfect samples form photon distribution and is very efficient as it scales as $\mathcal{O}{(\chi^{2})}$ for a properly right-orthogonalized MPS. This method can be especially useful in situations in which photon statistical behavior needs to be taken into account.
In this section, we want to approximate the continuous function $E(r,\theta)$ by a matrix product state $|\phi\rangle$, without executing the exponentially expensive operation that is exhaustively storing the function on a grid $L\times L$. To overcome this limitation, we then adopt a sampling-based approach exploiting the family of tensor-cross interpolations algorithms. In particular, we focus on the method described in Ritter et al. (2024); Núñez Fernández et al. (2025), which is based on the LU decomposition. The algorithm we employ takes the input, represented as a function, and generates an approximated tensor in the quantics MPS form. The final tensor is obtained by deterministically sampling the initial one, with a linear number of calls $\mathcal{O}(n\chi^{2})$, depending on the rank of the function.
@@
Figure 12: Run times of Eq. (10) for the three apertures showed in Fig. 10 compared to the FFT computation in green.
@@
It can be argued that since the Zernike polynomials also live on the aperture, the maximum bond dimension is somehow dominated by the discontinuity, regardless of the number of modes considered. The numerical experiment (see Fig. 12) shows run times of Eq. (10) for the three apertures shown earlier. Times are measured for several grid sizes defined by the number of qubits and are compared to a standard FFT calculation. In logarithmic time, it indicates that the run times for the MPS/MPO formulation are (up to log corrections) constant, with a constant depending on the complexity of $|\psi\rangle$, while the FFT computation scales linearly with system size.
@@
For an extended scene $I_{\text{obj}}$, the image formed in the focal plane is given by the convolution of the scene with the point spread function. Mathematically, this is expressed as $I_{\text{img}}=I_{\text{obj}}\ast\text{PSF}$. In the Fourier domain, this convolution simplifies to a multiplication $\hat{I}_{\text{img}}(f_{x},f_{y})=\hat{I}_{\text{obj}}(f_{x},f_{y})\cdot\hat{\text{PSF}}(f_{x},f_{y})$, where $\hat{I}$ and $\hat{\text{PSF}}$ are the Fourier transforms of the object and PSF, respectively. The function $\hat{\text{PSF}}$ is also known as the optical transfer function, which characterizes how different spatial frequencies are transmitted by the optical system. The extended scene $I_{\text{obj}}$ can also be seen as a state $|S\rangle$ such that the final image state reads
$I_{\text{img}}=I_{\text{obj}}\ast\text{PSF}\rightarrow\hat{U}_{\mathrm{iqft}}(\hat{U}_{\mathrm{qft}}|\mathrm{PSF}\rangle\times\hat{U}_{\mathrm{qft}}|\mathrm{S}\rangle).$ (14)
The algorithm used to perform the calculation of element-wise multiplication scales as $\mathcal{O}{(d\chi^{3})}$, a faster algorithm Michailidis et al. (2024) could be implemented to speed things up. Evaluating this process for a finite number of qubits $n=2\log_{2}(L)$ is exactly equivalent to the exact calculation on a finite grid of size $L\times L$ by matrix multiplication and fast Fourier transform (FFT). An implementation is showed in the plot below (see Fig. 13). We plot the run times versus a classical FFT implementation for comparisons for increasing approximation error $\epsilon$. The difference in scaling is similar to the simpler case of the PSF computation.
@@
Figure 13: Run times of the matrix product image formation for increasing precisions versus the standard FFT implementation. The MPS/MPO construction is showed on the bottom right and the circular aperture is used.
@@
In this formalism, the entire image formation chain can be computed efficiently, with varying levels of complexity depending on the source image, aperture, and aberrations. In addition, other components of realistic systems—such as polarization, detector modulation functions, and even noise—can also be incorporated. As the name suggests, Fourier optics is heavily based on the fact that classical optics becomes computationally faster in Fourier space as convolutions become multiplications. In the following section we will exploit this feature in more depth and apply it to other optical propagations.
V Wave optics and quantum mechanics

### v2

@@ IV Matrix Product States and Tensor Tree Networks
Nevertheless, it is clear that the error goes to zero rather differently due to the maximal number of parameters allowed by the SVD in these kinds of networks. We can show that a large $n$, a maximally random tensor tree has at most $2^{n}$ parameters. We take the example of the boundary binary tree (see right Fig. 10), the maximal number of parameters is $D_{\mathrm{max}}(n)=n(2\times 2)+\sum_{k=1}^{\log_{2}(n)-2}\frac{n}{2^{k}}\left(2^{2^{k+1}-2}\times 2^{2^{k+1}-1}\right)+2\times 2^{n-2}$. The first term $n(2\times 2)$ corresponds to the $n$ outside branch tensors and the last term corresponds to the two central tensors. The largest term inside the sum is $2^{n}-1$ so $D_{\mathrm{max}}(n)=4n+2^{n-1}+2.2^{n-2}=2^{n}+4n$, implying a value (up to a tiny $\mathcal{O}(n)$ term) of $\xi_{\infty}^{\mathrm{TTN}}=1$ for large $n$. The exponential term comes from top layers, required to encode maximal entanglement and the linear term $4n$ comes from leaves and small subtrees. The noise is treated as scale-separated, since it lives un in the tree.
A similar counting operation can be done for bulk tensor trees, and would have lead to the same result. It implies that in the thermodynamic limit (in contrast to matrix product states), tensor trees always compress the image, $\xi^{\mathrm{TTN}}_{\infty}\leq 1$ regardless of the level of approximation (and encodings) which is already an interesting fact on its own and will be further explain in the following. Therefore, tensor trees offer greater compressive power in the intermediate regime (see Fig. 11), as the number of parameters approaches the compression limit with increasing system size. This is illustrated in the table below, where tensor trees clearly outperform linear networks and continue to compress the image even for very small reconstruction error [^9] We do not show the results of FRQI on trees as they perform poorly compared to quantics..
@@
Table 1: Compression ratio for several errors with $L=2048$. The three firsts rows use FRQI with respectively a random walk (SARW), snake-like (S) and a Hilbert curve (H) path encoding. The fourth is the quantics tensor train $q\mathrm{TT}$ and the last is a quantics encoding on bulk tensor trees $q\mathrm{TTN}_{B}$.
In the case of trees, entanglement is distributed across a hierarchy of scales rather than along a single ordered chain, and is progressively pushed toward higher levels of the tree as reconstruction accuracy increases. When encoding an image using the quantics indexing, the resulting tensor indices are naturally ordered by spatial scale, from coarse to fine resolution. While this multi-scale indexing already improves the entanglement structure compared to a path encodings, the topology of the tensor network remains the main factor for compression. In a MPS, the chain topology forces noise-induced entanglement to be distributed uniformly across all links, leading to large bond dimensions at all scales.
By contrast, a binary TTN aligns naturally with the hierarchical structure induced by the quantics representation: local tensors capture fine-scale correlations, while global correlations and noise entropy are progressively pushed toward the upper layers of the tree. As a result, Haar entanglement appears only at the highest levels of the TTN, allowing strong and aggressive truncation of top bonds while preserving local image features. This scale separation of entropy explains why TTN representations provide more efficient compression of noisy images than MPS, even when both use the same quantics encoding. Since TTN is strictly more efficient in the noisy regime while remaining equivalent to MPS in the area-law limit, continuity of entanglement scaling implies that TTN is generically advantageous in the intermediate regime where weak violations of the area law are present, as is typical for noisy or high-precision data. When the entanglement structure becomes genuinely 2dd or exhibits scale-invariant behavior, PEPS and MERA offer more natural and efficient descriptions by adapting the tensor network geometry with the underlying entropy structure of the data.
Although comparisons with standard image compression algorithms are not showed here, we mention that tensor networks might have a big advantage at very large size and relative high error. Classical block-based compression schemes such as JPEG operate by partitioning the image into fixed-size blocks and applying local transforms within each block, thereby imposing a hard cutoff on correlations beyond the block scale. While effective for locally smooth images, this rigid locality prevents an adaptive representation of long-range or multi-scale correlations and leads to well-known blocking artifacts. In contrast, tensor network representations do not enforce fixed spatial partitions but instead truncate correlations adaptively across scales, allowing global structure to be preserved while suppressing noise. A comprehensive benchmark, although beyond the scope of this paper, on the performance of tensor networks across large datasets of diverse images would be necessary to rigorously demonstrate their relative strengths and weaknesses.
Despite these promising features, one of the main strength of tensor networks lies not only in data reduction itself but in their ability to perform complex operations efficiently. In particular, there are particularly suited for image transformations tasks, where operations can be applied directly at the level of the tensor network. The last section of this work will therefore be directed towards optical applications that could be significantly sped up using this formalism.
V Image processing and optical image formation
@@
Figure 12: Operations on images as MPS/MPO contraction on the left and TTN/TTO contraction on the right.
@@
Equivalently, it can be written as a unitary operator $\hat{U}$ on the space of qubits such that [^10] In this paper, we choose the quantum computation notation $\{\hat{X},\hat{Y},\hat{Z}\}$ instead of $\{\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z}\}$ for the Pauli matrices.
@@
In the following, we shall see how to incorporate our previous results on image encoding in the framework of a realistic optical system. The simplest of calculations is to compute the point spread function (PSF) of a classical optical system that contains phase aberrations. An image observed with that aberrated optical system will then be modeled by the convolution of the input image and the PSF. Let us start be defining a monochromatic coherent field as the plane wave
$E_{0}(r,\theta)=A(r,\theta)\mathrm{e}^{-ik\Phi(r,\theta)},$ (11)
@@
$E_{\mathrm{f}}(u,v)=\int r\mathrm{d}r\int\mathrm{d}\theta E_{0}(r,\theta)\mathrm{e}^{-ikr(u\cos\theta+v\sin\theta)},$ (12)
where $(u,v)$ are spatial coordinates in the focal plane, and $A(r,\theta)$ is the pupil function, typically a circular aperture that  [^11] We switch liberally from Cartesian to polar coordinates depending on cases.. For simplicity in the simulations, we choose the arbitrary function $\Phi(u,v)=\cos(u)\cos(u)+\cos(2u+5v)+\sin(uv)+\cos(2u+2v)^{13}$ and the apertures showed in Fig. 13. The point spread function is simply given by
$\text{PSF}(u,v)$ $=\left|E_{\mathrm{f}}(u,v)\right|^{2}$
$=|\mathcal{F}(E_{0})|^{2},$ (13)
which is the intensity of the field at the focal plane ($\mathcal{F}$ being the Fourier transform). In the following we will consider it as normalized $\text{PSF}(u,v)\rightarrow\text{PSF}(u,v)/|\text{PSF}(u,v)|$ such that it represents a photon probability distribution.
Figure 13: Three classic apertures considered for the simulations modulated by continuous phase field living on them. The middle case will be surprisingly easy to approximate with quantics TCI while the others will turn up to be more difficult.
Let us see how one defines this classical calculation in terms of MPS and MPO. For the time being, we admit that we successfully constructed a normalized MPS representation of $E_{0}$ that we note $|\phi\rangle$. The classical Fraunhofer propagation of the continuous field $E_{0}$ can be seen as a unitary propagation of a normalized quantum state $|\phi\rangle$ living in the Hilbert space $\mathcal{H}$ of infinitely many qubits. The non-unitary operation that computes the PSF of the optical system is, in this context, defined by the non-unitary process of quantum measurement. Mathematically, it boils down to the following dictionary
$E_{0}\leftrightarrow|\phi\rangle\in\mathcal{H},$
$\mathcal{F}_{xy}\{\bullet\}\leftrightarrow\hat{U}_{\mathrm{qft}}=\hat{U}^{(x)}_{\mathrm{qft}}\otimes\hat{U}^{(y)}_{\mathrm{qft}}$$\mathcal{H}\rightarrow\mathcal{H}$,
$E_{\mathrm{f}}\leftrightarrow|\varphi\rangle=\hat{U}_{\mathrm{qft}}|\phi\rangle\in\mathcal{H},$
$\mathrm{PSF}\leftrightarrow\|\mathrm{PSF}\rangle\rangle=|\varphi\rangle\odot|\varphi\rangle^{{\dagger}}\in\mathcal{H}_{\mathrm{cl}},$ (14)
where we use $\|.\rangle\rangle$ to denote MPS representations of classical fields or probability distributions living on $\mathcal{H}_{\mathrm{cl}}$, as opposed to quantum state vectors in $\mathcal{H}$ and where $\odot$ is the local contraction of physical indices, corresponding in the continuum limit to point-wise multiplication in the measurement basis  [^12] The algorithm used to perform the calculation of element-wise multiplication scales as $\mathcal{O}{(d\chi^{3})}$, a faster algorithm Michailidis et al. (2024) could be implemented to speed things up.. The local contraction  [^13] In this notation, $|\varphi\rangle\odot|\varphi\rangle^{{\dagger}}$ stores the diagonal values of the density matrix $\rho$ such that $\mathcal{H}_{\mathrm{cl}}\simeq\{\operatorname{diag}(\rho):\rho\in\mathcal{B}(\mathcal{H})\}.$ is the MPS analog of projective measurement, mapping a quantum state to a classical probability $\mathrm{PSF}=|\varphi|^{2}$, with normalization $\langle\varphi|\varphi\rangle=1$ (equivalently $\|\mathrm{PSF}\|_{L^{1}}=1$). This normalized PSF then can be seen a the spatial probability of photon events $\mathbb{P}[\bullet$,(u,v)],(u,v)] where $\bullet$ is the photon event. A summary of the discretized calculation steps is illustrated in Fig. 14 141414mention that a $\mathrm{FFTSHIFT}$ operation..).
@@
Let us mention, that the point-wise product operation is only needed if we absolutely require the PSF to be in a MPS format, to be applied later on to another MPO for example. But many operations can be done at the level of $|\varphi\rangle$. Direct contraction of the MPS $|\varphi\rangle$ followed by an absolute square operation can be done to plot quickly some parts of the large PSF for example. Also, contrary to the standard FFT calculation, the matrix product state formulation allows photon sampling without ever computing the photon statistical distribution (see Fig. 14) but by sampling the field $|\varphi\rangle$ after orthogonalization. The perfect sampling algorithm can be used to sample from this distribution Stoudenmire and White (2010); Ferris and Vidal (2012). As the name suggests, the algorithm draws perfect samples from photon distribution and is very efficient as it scales as $\mathcal{O}{(\chi^{2})}$ for a properly right-orthogonalized MPS. This method could be especially useful in situations in which photon statistical behavior needs to be taken into account, or to construct a joint distribution over photon events and aberrations in a bayesian fashion.
As said earlier, a low-rank matrix product operator construction of $\hat{U}_{\mathrm{qft}}$ has recently been developed and could be implemented here. To take advantage of this MPO formulation, the field $|\phi\rangle$ requires also a low-rank form. In this section, we proceed to approximate the continuous function $E_{0}(r,\theta)$ by a MPS, without executing the exponentially expensive operation that is exhaustively storing the function on a grid $L\times L$. To overcome this limitation, we then adopt a sampling-based approach exploiting the family of tensor-cross interpolations algorithms. In particular, we focus on the quantics tensor cross-interpolation (qqTCI) method described in Ritter et al. (2024); Núñez Fernández et al. (2025), which is based on the LU decomposition. The algorithm we employ takes the input, represented as a function, and generates an approximated MPS in the quantics interleaved representation. The final tensor is obtained by deterministically sampling the initial one, with a linear number of calls $\mathcal{O}(n\chi^{2})$, depending on the rank of the function.
@@
Figure 15: Run times of Eq. (V) for the three apertures showed in Fig. 13 compared to the fast Fourier transform (FFT) computation in green.
@@
It can be argued that since the phase function also live on the aperture, the maximum bond dimension is somehow dominated by the discontinuity, regardless of the degree of the polynomials considered. The numerical experiment (see Fig. 15) shows run times of Eq. (V) for the three apertures shown earlier, including time necessary to encode the field Eq. (11) with qqTCI. Times are measured for several grid sizes defined by the number of qubits and are compared to a standard FFT calculation  [^15] Package FFTW https://github.com/JuliaMath/FFTW.jl. In logarithmic time, it indicates that the run times for the MPS/MPO formulation are (up to log corrections) constant, with a constant depending on the complexity of $|\varphi\rangle$, while the FFT computation scales linearly with system size. In the figure above, we stop at around $n=25$ but size upwards of $n>50$ are possible since we are working with qqTCI and need not store an exponential number of values. This would not be the case in the next section since we will connect to the first section of this article and apply this formalism to the original image.
@@
For an extended scene $I_{\text{obj}}$, the image formed in the focal plane is given by the convolution of the scene with the point spread function. Mathematically, this is expressed as $I_{\text{img}}=I_{\text{obj}}\ast\text{PSF}$. In the Fourier domain, this convolution simplifies to a multiplication $\hat{I}_{\text{img}}(f_{x},f_{y})=\hat{I}_{\text{obj}}(f_{x},f_{y})\cdot\hat{\text{PSF}}(f_{x},f_{y})$, where $\hat{I}$ and $\hat{\text{PSF}}$ are the Fourier transforms of the object and PSF, respectively. The function $\hat{\text{PSF}}$ is also known as the optical transfer function, which characterizes how different spatial frequencies are transmitted by the optical system. As we saw earlier, the extended scene $I_{\text{obj}}$ can also be seen as a normalized quantum state MPS $|S\rangle$ such that the final image state reads
$I_{\text{img}}=I_{\text{obj}}\ast\text{PSF}\rightarrow\hat{U}_{\mathrm{iqft}}(\hat{U}_{\mathrm{qft}}||\mathrm{PSF}\rangle\rangle\odot\hat{U}_{\mathrm{qft}}|\mathrm{S}\rangle).$ (15)
Evaluating this tensor network operation for a finite number of qubits $n=2\log_{2}(L)$ is exactly equivalent to compute $I_{\mathrm{img}}$ finite grid of size $L\times L$ by matrix multiplication and FFT. An implementation is shown in the plot below (see Fig. 16) for the case of the circular aperture. We use as input the picture used in the beginning of this article, and put it through the optical system Eq. (15) for several sizes. We plot the run times versus a classical FFT implementation for comparisons for increasing approximation error $\epsilon$. Here we choose $\epsilon$ as the same cutoff in precision for both input MPS: the input image and the input field, while the QFT MPO and the multiplication run in machine precision. Also again, run times take into consideration the encodings of both MPS. The difference in scaling is similar to the simpler case of the PSF and the same conclusions can be drawn.
@@
Figure 16: Run times of the tensor network image formation for increasing precisions ($\epsilon_{1}=10^{-2}$$\epsilon_{2}=10^{-3}$$\epsilon_{3}=10^{-4}$) versus the standard FFT implementation. The MPS/MPO construction is shown on the bottom right and the circular aperture is used.
The dominant contribution in the entanglement of the entire chain comes from the input image approximation, but it might be useful to point out that, a time advantage can still remain even if the image MPS is no longer in the compressive regime $\xi>1$, since the rest of the contractions are very low rank. The entire computation can still be computationally efficient at large-scale even if the image MPS alone might be widely inefficient.
@@
In this formalism, the entire image formation chain can be computed efficiently, with varying levels of complexity depending on the source image, aperture, and aberrations. In addition, other components of realistic systems, such as polarization, detector modulation functions and even noise can also be incorporated. In particular, incoherent imaging and statistical optics are easily generalizable with these methods. As the name suggests, Fourier optics is heavily based on the fact that classical optics becomes computationally faster in Fourier space as convolutions become multiplications. In the following section we will exploit this feature in more depth and apply it to other classical optical propagations.
VI Wave optics and quantum mechanics